OpenStudy (anonymous):

((1/1+tanx))+((1/1+cotx))=1 Verify the Identity

OpenStudy (anonymous):

trig identities

OpenStudy (anonymous):

I know that but i'm having a hard time getting it to equal 1

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

((1/1+tanx))+((1/1+cotx))= 1/(cosx/sinx + sinx/sinx) + 1/(sinx/cosx + cosx/cosx)= 1/((cosx+sinx)/sinx) + 1/((sinx+cosx)/cosx) = sinx/(sinx+cosx) + cosx/(sinx+cosx) =(sinx+cosx) / (sinx+cosx) =1

OpenStudy (anonymous):

Thank you so much. I see what I did wrong now.

OpenStudy (anonymous):

ur welcome